Students sold peanut butter cookies for $6.50 a dozen. They sold chocolate chip cookies for $9.00 a dozen. They sold 85 dozen more peanut butter cookies than chocolate chip cookies. The total dollar amount sold was 4055.50. How many dozen peanut butter cookies were sold?
Answer by mananth(16946)

Students sold peanut butter cookies for $6.50 a dozen. They sold chocolate chip cookies for $9.00 a dozen. They sold 85 dozen more peanut butter cookies than chocolate chip cookies. The total dollar amount sold was 4055.50. How many dozen peanut butter cookies were sold?
Let number of chocolate chip cookies sold be x
peanut butter cookies sold = x+85.
..
peanut butter cookies for $6.50 a dozen.
chip cookies for $9.00 a dozen
..
Total receipts = 4055.50
9x+6.5(x+85)=4055.50
9x+6.5x+552.5 =4055.50
15.5x=4055.50-552.50
15.5x=3503
x=3503/15.5
x= 226 chocolate chip cookies
226+85 = 311 peanut butter cookies